OpenStudy (anonymous):

Solve the equation 2/3 (x − 4) = 2x for x

OpenStudy (anonymous):

(2/3)(x-4)=2x [(2x)/3]-(8/3) = 2x -(8/3) = 2x - [(2x)/3] -(8/3) = [(6x)/3]-[(2x)/3] -(8/3) = (4x)/3 -8/3 * (3/4) = x -24/12 = x -2 = x